OpenStudy (anonymous):

Is my answer right? Simplify the given expression. sq root -54 + sq root -96 Answer: - 5 sq root 6

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Is that it? Are you studying square roots of negative numbers?

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{-54} + \sqrt{-96} \neq -5\sqrt{6}\]

OpenStudy (anonymous):

there are gonna be some \(\large{i}\)s in there